# DeepDetect: Learning All-in-One Dense Keypoints

### v1

@@ III-B ESPNet based Training
In the training stage (see Figure 3), ESPNet is trained on the prepared dataset for $100$ epochs with a learning rate of $0.001$ and batch size of $64$, using cosine annealing scheduler for smoother convergence. We used Python and PyTorch [32] for coding. The size of input images is kept consistent at $3\times 480\times 480$ for the model and the predicted output mask is re-sized to match the spatial dimensions of the original image. Given a target mask $M(I)=y\in\{0,1\}^{H\times W}$ and model’s raw output $z\in\mathbb{R}^{H\times W}$, the BCE loss $L(y,z)$ is described as:
@@ 66 Correct Matches
Results show that DeepDetect achieves state-of-the-art performance. On the Oxford dataset [11], it gives an average repeatability score of 0.9582, surpassing other keypoint detectors by a significant margin. The results for average keypoint density are presented in Table I, where DeepDetect achieves the highest average density of 0.5143. Table I also reports average foreground-to-keypoint ratio on $18$ custom-selected images with ground-truth binary masks. DeepDetect achieves good ratios, indicating that its keypoints are not only dense but also well-concentrated in semantically important regions in the images. With default thresholds, the classical detectors yield higher F-KP ratios but significantly lower keypoint densities. Conversely, reducing thresholds to increase density decreases the F-KP ratio, highlighting the presence of noisy keypoints. For instance, FAST and AGAST with low thresholds, achieve high densities with low F-KP ratios, showing their inability to concentrate on salient regions. DeepDetect produces considerably higher densities while maintaining strong foreground focus. Visual comparisons in Figure 2 clearly show that DeepDetect focuses on semantically meaningful regions and provides very high keypoint density. Table II and Figure 6 show that DeepDetect provides substantially higher correct matches and consistently outperforms classical detectors. The strength of DeepDetect lies in its ability to combine the geometric robustness of keypoint detectors, semantic awareness of edge detectors, and adaptability to extreme image variations gained through the use of supervision masks generated with different thresholds for normal and challenging scenes. DeepDetect demonstrates stronger generalization across degraded visual conditions due to its intelligent nature.
@@ Conclusion
We introduce DeepDetect, an intelligent, all-in-one, dense keypoint detector that unifies the robustness of 7 keypoint detectors with the semantic awareness of 2 edge detectors using deep learning. By leveraging multi-detector and multi-edge fusion, we constructed diverse supervision masks that enable DeepDetect to learn rich, dense, and semantically focused keypoints in the images. Extensive experiments on Oxford Affine Covariant Regions demonstrate that DeepDetect consistently outperforms other detectors in repeatability, keypoint density, and correct matches while maintaining robustness under extreme image variations. Its ability to produce highly dense, repeatable, and semantically meaningful keypoints without manual tuning, make it a well suited detector for dense 3D reconstruction, visual SLAM, AR/VR, object tracking, and autonomous navigation applications in adverse environments.

### v2

@@ III-B ESPNet based Training
In the training stage (see Figure 3), ESPNet is trained on the prepared dataset for $100$ epochs with a learning rate of $0.001$ and batch size of $64$, using cosine annealing scheduler for smoother convergence. We used Python, OpenCV, and PyTorch [32] for coding. The input image size is kept consistent at $3\times 480\times 480$ for the model and its predicted output mask is re-sized to match the spatial dimensions of the original image. Given a target mask $M(I)=y\in\{0,1\}^{H\times W}$ and model’s raw output $z\in\mathbb{R}^{H\times W}$, the BCE loss $L(y,z)$ is described as:
@@ 66 Correct Matches
Results show that DeepDetect achieves state-of-the-art performance. On the Oxford dataset [11], it gives an average repeatability score of 0.9582, surpassing other keypoint detectors by a significant margin. The results for average keypoint density are presented in Table I, where DeepDetect achieves the highest average density of 0.5143. Table I also reports average foreground-to-keypoint ratio on $18$ custom-selected images with ground-truth binary masks. DeepDetect achieves good ratios, indicating that its keypoints are not only dense but also well-concentrated in semantically important regions in the images. With default thresholds, the classical detectors yield higher F-KP ratios but significantly lower keypoint densities. Conversely, reducing thresholds to increase density decreases the F-KP ratio, highlighting the presence of noisy keypoints. For instance, FAST and AGAST with low thresholds, achieve high densities with low F-KP ratios, showing their inability to concentrate on salient regions. DeepDetect produces considerably higher densities while maintaining strong foreground focus. Visual comparisons in Figure 2 clearly show that DeepDetect focuses on semantically meaningful regions and provides very high keypoint density. Table II and Figure 6 show that DeepDetect provides substantially higher correct matches and consistently outperforms classical detectors. The strength of DeepDetect lies in its ability to combine the geometric robustness and diversity of keypoint and edge detectors, and adaptability to extreme image variations gained through the use of supervision masks generated with different thresholds for normal and challenging scenes. DeepDetect demonstrates stronger generalization across degraded visual conditions due to its intelligent nature.
@@ Conclusion
We introduce DeepDetect, an intelligent, all-in-one, dense keypoint detector that unifies the robustness and distinctiveness of 7 keypoint detectors and 2 edge detectors using deep learning. By leveraging multi-detector and multi-edge fusion, we constructed diverse supervision masks that enable DeepDetect to learn rich, dense, and semantically focused keypoints in the images. Extensive experiments on Oxford Affine Covariant Regions demonstrate that DeepDetect consistently outperforms other detectors in repeatability, keypoint density, and correct matches while maintaining robustness under extreme image variations. Its ability to produce highly dense, repeatable, and semantically meaningful keypoints without manual tuning, make it a well suited detector for dense 3D reconstruction, visual SLAM, AR/VR, object tracking, and autonomous navigation applications in adverse environments.